A random sample of 10 ​venture-capital investments in the fiber optics business sector yielded the following​ data, in millions of dollars. Determine a 95​% confidence interval for the mean​ amount, mu​, of all​ venture-capital investments in the fiber optics business sector. Assume that the population standard deviation is ​$2.43 million.​ (Note: The sum of the data is ​$62.59 ​million.)

The confidence interval for mu is from ​$ _____million to ​$______million. ​(Round to two decimal places as​ needed.)
Answer #1

Answer

sample size is n = 10

sum of data is 62.59

so, mean = (sum)/(sample size) = 62.59/10 = 6.259

population standard deviation is sigma = 2.43

z score for 95% confidence interval is 1.96 (using z table)

Confidence interval =

The confidence interval for mu is from ​$4.75 million to ​$7.77 million